OpenStudy (anonymous):

Which equations represent the line shown in the graph? Choose all answers that are correct The ploted points that make the line are (1,5) and (-4,-3) A. y+4=8/5(x+3) B. -8x+5y=17 C. y=8/5x+17/5 D. 5x-8y=17 Please explain your answer and thank you for helping!

OpenStudy (anonymous):

haha sup girl xD equation: y-y1=m(x-x1) but find the slope: m=y2-y1/x2-x1 x1,y2 are the point

OpenStudy (anonymous):

I need graph

OpenStudy (anonymous):

Nvm question finally fully loaded

OpenStudy (anonymous):

lmao xD wow

OpenStudy (anonymous):

@atreyu6s3x6 You again, hi!! XD Thanks for helping!!

OpenStudy (anonymous):

yes.......... it is me...................<.< lol no problem buddy xD